Question 12 (0.15 points) A companys board of directors approved a 3-for-1 stock split. If you owned 70 shares before the st please answer 12 & 13. thank you.
Answer #1

Answer:

Question 12)

Stock split =3 to1

Number of shares holds N=70

Current price of share P=23.47/Share

So new share price =P/3=23.47/3=$7.823

Question 13)

Stock split =2 to1

Number of shares holds N=100

Current price of share P=11.88/Share

So new share price =P/2=11.88/2=$5.94